# Complete moduli of Enriques surfaces in Horikawa’s model

### 06 Dec 2023 v1

@@ 1. Introduction
Enriques surfaces are quotients of K3 surfaces by basepoint free involutions. They satisfy $2K\sim 0$ and $q=0$ and occupy a place somewhere in between rational and K3 surfaces. Unlike K3 surfaces, there are only finitely many moduli spaces of Enriques surfaces, see [GH16]. Each of them parameterizes the same surfaces, with some finite data attached. There are many classical constructions due to Enriques, Castelnuovo, Fano, etc, see [CD89] for more details. The model we use here is due to Horikawa [Hor78], see also [BPVdV84, Sec.23]. It is also known as the moduli space $F_{\rm En,2}$ of Enriques surfaces with a numerical polarization of degree $2$.
Consider a surface $Y={\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ with an involution $\tau\colon(x,y)\to(-x,-y)$, and let $B\in|-2K_{Y}|$ be a smooth $\tau$-invariant divisor not passing through the four points with $x,y\in\{0,\infty\}$ fixed by $\tau$. The double cover $\pi\colon X\to Y$ branched in $B$ is a K3 surface with a del Pezzo involution $\iota_{{\rm dP}}$ such that $Y=X/\iota_{{\rm dP}}$. The involution $\tau$ on $Y$ lifts to a basepoint-free Enriques involution $\iota_{{\rm En}}$ on $X$ commuting with $\iota_{{\rm dP}}$ and the quotient $Z=X/\iota_{{\rm En}}$ is an Enriques surface. We have a commutative diagram
[…]
The pullback ${\mathcal{L}}=\pi^{*}({\mathcal{L}}_{Y})$ of the ample degree $2$ line bundle ${\mathcal{L}}_{Y}={\mathcal{O}}(1,1)$ on the quadric $Y$ is a polarization on $X$ of degree $4$. It is also the pullback of a numerical polarization $[{\mathcal{L}}_{Z}]\in(\operatorname{Pic}Z)/{\mathbb{Z}}_{2}$ of degree $2$ via $\psi$.
Let $R$ be the ramification divisor of the quotient map $\pi\colon X\to Y$ by the del Pezzo involution. Then $\iota_{{\rm En}}(R)=R$ so we may define an integral divisor $R_{Z}:=\frac{1}{2}\psi_{*}(R)$. Alternatively, the involution $\iota_{{\rm dP}}$ on $X$ descends to an involution $\tau_{{\rm dP}}$ on $Z$, and $R_{Z}$ is the ramification divisor of the quotient map $\rho\colon Z\to W$ by $\tau_{{\rm dP}}$. Since $R=\psi{}^{*}(R_{Z})=\frac{1}{2}\pi^{*}(B)$ is an ample divisor, $R_{Z}$ is ample as well. Note that ${\mathcal{O}}(R_{Z})={\mathcal{L}}_{Z}^{\otimes 2}\in\operatorname{Pic}Z$ is a well-defined line bundle.
The moduli space $F_{\rm En,2}$ and its Baily-Borel compactification ${\overline{F}}_{\rm En,2}^{\rm BB}$ were studied by Sterk [Ste91]. In this paper we describe the modular compactification $F_{\rm En,2}\hookrightarrow{\overline{F}}_{\rm En,2}$ in the moduli space of KSBA-stable pairs $(Z,\epsilon R_{Z})$ for $0<\epsilon\ll 1$. See [Kol23] for their general theory. Our main result, in Section 5.2, is
@@
In Sections 6, 7 we describe all stable pairs parametrized by the boundary of ${\overline{F}}_{\rm En,2}$. In Section 4 we also provide a detailed description of various models of Enriques degenerations. Just as weighted graphs encode semistable degenerations of curves, $K$-trivial semistable (i.e. Kulikov) degenerations of K3 surfaces ${\mathcal{X}}\to(C,0)$ are encoded by integral-affine structures on the $2$-sphere, or $\operatorname{IAS^{2}}$ [GHK15a, Eng18, EF21].
An $\operatorname{IAS^{2}}$ is a collection of local embeddings of $S^{2}$ minus a finite set into the flat plane, which differ on overlaps by $\operatorname{SL}_{2}({\mathbb{Z}})\ltimes{\mathbb{Z}}^{2}$. Complete triangulations of $\operatorname{IAS^{2}}$ which take their vertices in ${\mathbb{Z}}^{2}$, under the local embedding, describe the dual complexes $\Gamma({\mathcal{X}}_{0})$ of Kulikov degenerations. In this paper, we realize diagram (1) on the integral-affine level, by constructing $\Gamma({\mathcal{X}}_{0})$ together with two commuting involutions $\iota_{{\rm En},\,{\rm IA}}$ and $\iota_{{\rm dP},{\rm IA}}$ of the $\operatorname{IAS^{2}}$.
@@
The plan of the paper is as follows: In Section 2 we further discuss Horikawa’s construction and recall the description of the moduli space $F_{\rm En,2}$ after Sterk [Ste91]. We also explain how it relates to the moduli space of (ordinary, not numerically) polarized degree $2$ Enriques surfaces as in [Vie95, GH16]. We describe morphisms between $F_{\rm En,2}$, $F_{(2,2,0)}$ and the moduli $F_{(10,10,0)}$ of unpolarized Enriques surfaces. Finally, we briefly recall the theory of KSBA stable pairs and their compact moduli.
In Section 3 we recall the cusp diagrams of $F_{(2,2,0)}$, $F_{(10,10,0)}$, $F_{\rm En,2}$ and determine how they map to each other. Next, we describe Coxeter diagrams associated to each of the five 00-cusps and the $18$ $1$-cusps of $F_{\rm En,2}$. The diagrams are the same as in Sterk [Ste91] but we find them in a novel way, “folding by involutions” the Coxeter diagrams of the lattices $U\oplus E_{8}^{2}$ and $U(2)\oplus E_{8}^{2}$ corresponding to the two 00-cusps of $F_{(2,2,0)}$. This is the combinatorial heart of the paper. An idea from [AET23, AE22] employed here is that one can read off degenerations of K3 surfaces directly from Coxeter diagrams. Consequently, we are able to read off degenerations of Enriques surfaces from the Coxeter diagrams of $F_{\rm En,2}$.
@@ 2. General setup
2.1. Horikawa’s construction
We begin by exploring diagram (1) in the introduction in more detail. The surface $Y={\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ is toric and the line bundle ${\mathcal{O}}(4,4)$ has, as its polytope, a square $Q$ with sides of lattice length $4$, see Figure 1. The surface $W$ is toric as well, for the same polytope but for the even sublattice ${\mathbb{Z}}^{2}_{{\rm ev}}=\{(a,b)\mid a+b\in 2{\mathbb{Z}}\}$ shown by gray dots. Vectors in ${\mathbb{Z}}^{2}_{{\rm ev}}$ are in bijection with monomials $x^{a}y^{b}$ invariant under the involution $\tau\colon(x,y)\to(-x,-y)$. Here, we freely identify monomials $1,x,\dotsc x^{4}$ with $x_{0}^{4},x_{0}^{3}x_{1},\dotsc,x_{1}^{4}$, and $1,y,\dotsc,y^{4}$ with $y_{0}^{4},y_{0}^{3}y_{1},\dotsc,y_{1}^{4}$.
Figure 1. Polytope $Q$ and lattices for the toric surfaces $Y$ and $W$
@@ 2.1. Horikawa’s construction
which together with the identity form a Klein-four group. Both $\iota_{{\rm En}}$ and $\iota_{{\rm Nik}}$ are lifts of $\tau$. A nonvanishing $2$-form on $X$ is
@@
Horikawa [Hor78] analyzed in some detail the sets of possible equations $f(x,y)$ and the maps from various opens subsets of ${\mathbb{P}}^{12}$ to the period domain ${\mathbb{D}}/\Gamma$ and its Baily-Borel compactification, introduced in the next section. In particular, he showed that certain mildly singular $f(x,y)$ vanishing at a torus-fixed point correspond to Coble surfaces, which are $S_{2}$-quotients of nodal K3 surfaces.
@@ 2.2. Period domains
2.3. Relation to polarized Enriques surfaces
Another perspective is provided by the moduli space of Enriques surfaces with a degree $2$ polarization. By Viehweg’s theory [Vie95], there exists a quasi-projective coarse moduli space $\widehat{F}_{\rm En,2}$ parameterizing them. By [GH16, Proposition 4.1], we have a degree $2$tale morphism
$\widehat{F}_{\rm En,2}\rightarrow F_{\rm En,2}.$
This can be understood geometrically as follows. If $[\mathcal{L}]$ denotes the numerical equivalence class of an ample line bundle on an Enriques surface $S$, then the fiber of $(S,[\mathcal{L}])$ consists of $(S,\mathcal{L})$ and $(S,\mathcal{L}\otimes\omega_{S})$. By [Knu20], $\widehat{F}_{\rm En,2}$ is irreducible.
Remark 2.8.
Given a compactification $F_{\rm En,2}\subset\overline{F}_{\rm En,2}$, there is an induced compactification of $\widehat{F}_{\rm En,2}$ by taking the normalization of $\overline{F}_{\rm En,2}$ in the field of rational functions $\mathbb{C}(\widehat{F}_{\rm En,2})$.
@@ 3.3. Cusp diagram of $F_{\rm En,2}$
Sterk [Ste91] computed the cusp diagram for ${\overline{F}}_{\rm En,2}^{\rm BB}$. There are five 00-cusps for which we use Sterk’s numbering $1,2,3,4,5$. There are also $18$ distinct $1$-cusps.
@@ 3.7. Folding Coxeter diagrams by involutions
where $U_{e}$ is the unipotent subgroup, which acts trivially on ${\overline{T}}_{\rm En}$. We define $\Gamma_{{\rm dP},e}$ and $\Gamma_{{\rm En},e}$ similarly.
Denote by $W(\Gamma_{{\rm dP},e})$ the reflection subgroup of $\Gamma_{{\rm dP},e}$. Its Coxeter diagram $G_{{\rm dP},e}$ is one of the two Coxeter diagrams in Figure 5. Denote by $W(\Gamma_{{\rm En},2,e})$ the reflection subgroup of $\Gamma_{{\rm En},2,e}$; it is generated by reflections in the roots $\alpha\in T_{\rm En}$ with $\alpha\cdot e=0$.
@@ 5.1. Toroidal compactification for the Coxeter fans
The Coxeter semifan ${\mathfrak{F}}_{\rm cox}$ is the fan with support $\overline{{\mathcal{C}}}_{\mathbb{Q}}$ whose maximal cones are chambers of $W$, i.e. ${\mathfrak{C}}$ and its $W$-images.
@@
    Type III cones. These are in bijection with the elliptic subdiagrams of $G$.
@@
By Lemma 2.6 there is an immersion $j\colon F_{\rm En,2}\to F_{(2,2,0)}$ whose image is a Noether-Lefschetz locus in $F_{(2,2,0)}$. The normalization of $j(F_{\rm En,2})$ is then a toroidal compactification ${\overline{F}}_{\rm En,2}^{\rm cox}$ for the collection of fans $\{{\mathfrak{F}}^{k}_{\rm cox}\}_{k=1,2,3,4,5}$, one for each of 00-cusp of $F_{\rm En,2}$. The fans ${\mathfrak{F}}^{k}_{\rm cox}$ are the intersections of the above fans ${\mathfrak{F}}_{r}$ in the lattices ${\overline{T}}_{\rm dP}=(18,0,0)_{1}$ and $(18,2,0)_{1}$ with the sublattices ${\overline{T}}_{\rm En}=(10,10,0)_{1}$ and $(10,8,0)_{1}$, as in Section 3. By Lemma 3.8 and Corollary 3.12 these five fans are the Coxeter fans for the folded Coxeter diagrams $G_{r}^{k}$ of Figures 7, 8. By [Ste91] the induced groups acting on $e^{\perp}/e$ are of the form $\Gamma_{k}=\operatorname{Aut}(G_{r}^{k})\ltimes W(G_{r}^{k})$
@@
For $k=1,2,3,4,5$, the numbers of Type II + Type III rays in ${\mathfrak{F}}^{k}_{\rm cox}/\Gamma_{k}$ are $4+4$, $2+8$, $3+15$, $4+12$, $5+17$. The toroidal compactification ${\overline{F}}_{\rm En,2}^{\rm cox}$ has $10+56=66$ Type II + Type III divisors.

### 28 Dec 2023 v2

# Compact moduli of Enriques surfaces with a numerical polarization of degree 2

@@ 1. Introduction
Enriques surfaces are quotients of K3 surfaces by basepoint free involutions. They satisfy $2K\sim 0$ and $q=0$ and occupy a place somewhere in between rational and K3 surfaces. Unlike K3 surfaces, there are only finitely many moduli spaces of polarized Enriques surfaces, see [GH16]. Each of them parameterizes the same surfaces, with some finite data attached.
In this paper we consider the moduli space $F_{\rm En,2}$ of pairs $(Z,[{\mathcal{L}}_{Z}])$, where $Z$ is an Enriques surface with $ADE$ singularities and $[{\mathcal{L}}_{Z}]\in\operatorname{Pic}Z/{\mathbb{Z}}_{2}$ is an ample numerical polarization of degree $2$. Equivalently, this is the moduli space of $ADE$ Enriques surfaces $(Z,{\mathcal{M}})$ with a $2$-divisible polarization ${\mathcal{M}}={\mathcal{L}}_{Z}^{\otimes 2}\in\operatorname{Pic}Z$ of degree $8$. The Baily-Borel compactification ${\overline{F}}_{\rm En,2}^{\rm BB}$ of this space was described by Sterk [Ste91].
By the classification of big and nef linear system on Enriques surfaces [Cos83], see also [CD89, CDL24], the linear system $|{\mathcal{M}}|$ is basepoint free and defines a double cover $\rho\colon Z\to W$ to a quartic del Pezzo surface with $4A_{1}$ or $A_{3}+2A_{1}$ singularities. The ramification divisor $R_{Z}\in|{\mathcal{M}}|$ of $\rho$ is ample and the pair $(Z,\epsilon R_{Z})$ is log canonical for any $0<\epsilon\ll 1$. Thus, the moduli space $F_{\rm En,2}$ admits a geometric, modular compactification ${\overline{F}}_{\rm En,2}$ in the KSBA moduli space of pairs $(Z,\epsilon R_{Z})$ with semi log canonical singularities, $K_{Z}\equiv 0$ and ample ${\mathbb{Q}}$-Cartier divisor $R_{Z}$. See [Kol23] for their general theory. Our main result, in Section 5.2, is
@@
In Sections 6, 7 we describe all stable pairs parametrized by the boundary of ${\overline{F}}_{\rm En,2}$. The irreducible components of these pairs turn out to be surfaces that naturally correspond to the ABCDE Dynkin diagrams. They generalize the ADE surfaces of [AT21] that appear on the boundary of K3 moduli spaces.
In Section 4 we also provide a detailed description of some nice models of Enriques degenerations, which are slightly more singular than simple normal crossing. Instead, they are dlt (divisorially log terminal, see [KM98, Def. 2.37]).
Just as weighted graphs encode semistable degenerations of curves, $K$-trivial semistable (i.e. Kulikov) degenerations of K3 surfaces ${\mathcal{X}}\to(C,0)$ are encoded by integral-affine structures on the $2$-sphere, or $\operatorname{IAS^{2}}$ [GHK15a, Eng18, EF21]. An $\operatorname{IAS^{2}}$ is a collection of local embeddings of $S^{2}$ minus a finite set into the flat plane, which differ on overlaps by $\operatorname{SL}_{2}({\mathbb{Z}})\ltimes{\mathbb{Z}}^{2}$. Complete triangulations of $\operatorname{IAS^{2}}$ which take their vertices in ${\mathbb{Z}}^{2}$, under the local embedding, describe the dual complexes $\Gamma({\mathcal{X}}_{0})$ of Kulikov degenerations. In this paper, we realize Diagram (1) on the integral-affine level, by constructing $\Gamma({\mathcal{X}}_{0})$ together with two commuting involutions $\iota_{{\rm En},\,{\rm IA}}$ and $\iota_{{\rm dP},{\rm IA}}$ of the $\operatorname{IAS^{2}}$.
@@
The plan of the paper is as follows. In Section 2 we discuss the model of Enriques surfaces which we use in this paper. We also recall the description of the moduli space $F_{\rm En,2}$ after Sterk [Ste91]. Then we describe morphisms between $F_{\rm En,2}$, $F_{(2,2,0)}$ and the moduli $F_{(10,10,0)}$ of unpolarized Enriques surfaces. Finally, we briefly recall the theory of KSBA stable pairs and their compact moduli.
In Section 3 we recall the cusp diagrams of $F_{(2,2,0)}$, $F_{(10,10,0)}$, $F_{\rm En,2}$ and determine how they map to each other. Next, we describe Coxeter diagrams associated to each of the five 00-cusps and the $9$ $1$-cusps of $F_{\rm En,2}$. The diagrams are the same as in Sterk [Ste91] but we find them in a different way, “folding by involutions” the Coxeter diagrams of the lattices $U\oplus E_{8}^{2}$ and $U(2)\oplus E_{8}^{2}$ corresponding to the two 00-cusps of $F_{(2,2,0)}$. This is the combinatorial heart of the paper. An idea from [AET23, AE22] employed here is that one can read off degenerations of K3 surfaces directly from Coxeter diagrams. Consequently, we are able to read off degenerations of Enriques surfaces from the Coxeter diagrams of $F_{\rm En,2}$.
@@
We work over the complex numbers, although most of the results can be generalized to any field of characteristic $\neq 2$.
@@ 2. General setup
2.1. The main diagram, general case
Consider a surface $Y={\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ with an involution $\tau\colon(x,y)\to(-x,-y)$ and quotient $W=Y/\tau$. Let $B\in|-2K_{Y}|$ be a $\tau$-invariant divisor with at worst $ADE$ singularities not passing through the four points with $x,y\in\{0,\infty\}$ fixed by $\tau$. The double cover $\pi\colon X\to Y$ branched in $B$ is a K3 surface with a del Pezzo involution $\iota_{{\rm dP}}$ so that $Y=X/\iota_{{\rm dP}}$. The involution $\tau$ on $Y$ lifts to a basepoint-free Enriques involution $\iota_{{\rm En}}$ on $X$ commuting with $\iota_{{\rm dP}}$ and the quotient $Z=X/\iota_{{\rm En}}$ is an Enriques surface. The second lift of $\tau$ is a Nikulin involution and the quotient $Z^{\prime}=X/\iota_{{\rm Nik}}$ is a K3 surface with eight $A_{1}$ singularities and possibly more. This gives the following commutative diagram:
[…]
The surface $Y={\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ is toric and the line bundle ${\mathcal{O}}(4,4)$ has, as its polytope, a square $Q$ with sides of lattice length $4$, shown in the left panel of Figure 1. The surface $W$ is toric as well, for the same polytope but for the even sublattice ${\mathbb{Z}}^{2}_{{\rm ev}}=\{(a,b)\mid a+b\in 2{\mathbb{Z}}\}$ shown by gray dots. It is a quartic del Pezzo surface with four $A_{1}$ singularities. Vectors in ${\mathbb{Z}}^{2}_{{\rm ev}}$ are in bijection with monomials $x^{a}y^{b}$ invariant under the involution $\tau\colon(x,y)\to(-x,-y)$. Here, we freely identify monomials $1,x,\dotsc x^{4}$ with $x_{0}^{4},x_{0}^{3}x_{1},\dotsc,x_{1}^{4}$, and $1,y,\dotsc,y^{4}$ with $y_{0}^{4},y_{0}^{3}y_{1},\dotsc,y_{1}^{4}$.
Figure 1. Polytopes $Q$ and lattices for the toric surfaces $Y$ and $W$
@@ 2.1. The main diagram, general case
which together with the identity form a Klein-four group. Both $\iota_{{\rm En}}$ and $\iota_{{\rm Nik}}$ are lifts of $\tau$. On an affine subset of $X$ a nonvanishing $2$-form is given by
@@
Let $R$ be the ramification divisor of $\pi$. The involution $\iota_{{\rm dP}}$ on $X$ descends to an involution $\tau_{{\rm dP}}$ on $Z$, and $W=Z/\tau_{{\rm dP}}$. Let $R_{Z}$ and $B_{W}$ be the ramification and branch divisors of $\rho$. Then $R=\psi^{*}(R_{Z})$ and $R_{Z}=\frac{1}{2}\psi_{*}(R)$. Since $R=\frac{1}{2}\pi^{*}(B)$ is an ample divisor, $R_{Z}$ is ample as well. One has ${\mathcal{O}}(R_{Z})={\mathcal{L}}_{Z}^{\otimes 2}\in\operatorname{Pic}Z$.
Horikawa [Hor78b] analyzed in some detail the sets of possible equations $f(x,y)$ and the maps from various opens subsets of ${\mathbb{P}}^{12}$ to the period domain ${\mathbb{D}}/\Gamma$ and its Baily-Borel compactification, introduced in the next section. In particular, he showed that certain mildly singular $f(x,y)$ vanishing at a torus-fixed point correspond to Coble surfaces, which are $S_{2}$-quotients of nodal K3 surfaces.
@@
2.2. The main diagram, special case
The previous section describes the general case, when the K3 cover $X$ is non-unigonal. The special case corresponds to a Heegner divisor in $F_{\rm En,2}$ for which $(Y,L)=({\mathbb{P}}(1,1,2),{\mathcal{O}}(4))$ is a singular quadric. The toric surfaces $Y$ and $W$ correspond to the same polytope $Q$ shown in the right panel of Figure 1 but for different lattices: ${\mathbb{Z}}^{2}$ and ${\mathbb{Z}}^{2}_{{\rm ev}}$. The surface $W$ is a quartic del Pezzo surface with $A_{3}+2A_{1}$ singularities.
There are still $13$ even monomials giving a family over $U\subset{\mathbb{P}}^{12}$. However, in this case $\operatorname{Aut}(W)$, the centralizer of $\tau$ in $\operatorname{Aut}(Y)$, is three-dimensional, equal to $S_{2}\ltimes({\mathbb{C}}^{*}{}^{2}\ltimes{\mathbb{C}})$. So there is only a $9$-dimensional family of non-isomorphic surfaces.
Remark 2.1.
To our knowledge, Diagram (1) minus $Z^{\prime}$ first appeared in [Cos83, 6.3.1], see also [CD89, Cor. 4.7.2]. Horikawa model [Hor78a] is a birationally isomorphic version of this diagram. Ultimately, it can be traced back to the Enriques’ double plane model [Enr06]. We refer the reader to [CDL24] for a detailed historical account and many other projective models of Enriques surfaces.
Remark 2.2.
The entire Diagram (1) is intrinsic to the pair $(Z,[{\mathcal{L}}_{Z}])$, in both the general and special cases. Indeed, $W=\varphi_{|{\mathcal{L}}_{Z}^{\otimes 2}|}(Z)$, $Y=\operatorname{Spec}{\mathcal{O}}_{W}\otimes{\mathcal{O}}_{W}(A)$ where $A$ is the generator of $\operatorname{Tors}\operatorname{Cl}(W)={\mathbb{Z}}_{2}$, and $X=Y\times_{W}Z$. One has
$Z^{\prime}=\operatorname{Spec}{\mathcal{O}}_{W}\oplus{\mathcal{O}}_{W}(K_{W})\quad\text{and}\quad Z=\operatorname{Spec}{\mathcal{O}}_{W}\oplus{\mathcal{O}}_{W}(K_{W}+A)$
with the multiplications defined by the divisor $B_{W}\in|-2K_{W}|=|-2(K_{W}+A)|$.
@@ 3.3. Cusp diagram of $F_{\rm En,2}$
Sterk [Ste91] computed the cusp diagram for ${\overline{F}}_{\rm En,2}^{\rm BB}$. There are five 00-cusps for which we use Sterk’s numbering $1,2,3,4,5$. There are also $9$ distinct $1$-cusps.
@@ 3.7. Folding Coxeter diagrams by involutions
where $U_{e}$ is the unipotent subgroup, which acts trivially on $e^{\perp}/e={\overline{T}}_{\rm En}$. We define $\Gamma_{{\rm dP},e}$ and $\Gamma_{{\rm En},e}$ similarly.
Denote by $W(\Gamma_{{\rm dP},e})$ the reflection subgroup of $\Gamma_{{\rm dP},e}$. Its Coxeter diagram $G({\overline{T}}_{{\rm dP}})$ is one of the two Coxeter diagrams in Figure 5. Denote by $W(\Gamma_{{\rm En},2,e})$ the reflection subgroup of $\Gamma_{{\rm En},2,e}$; it is generated by reflections in the roots $\alpha\in T_{\rm En}$ with $\alpha\cdot e=0$.
@@ 3.9. $1$-cusps of $F_{\rm En,2}$ by folding
Remark 3.15.
The idea that folded diagrams may be relevant to compactifying $F_{\rm En,2}$ implicitly appears in [Ste91], e.g. there is a folded $A_{15}$ diagram in Fig. 16. We found that once the K3 case understood, the folding, when applied to the correct space—which is $F_{(2,2,0)}$ and not $F_{4}$—completely solves the Enriques case. Note that the moduli space $F_{4}$ of quartic K3 surfaces has a unique 00-cusp with a non-reflective hyperbolic lattice, but the moduli space $F_{(2,2,0)}$ of hyperelliptic K3 surfaces of degree $4$ has two 00-cusps with reflective hyperbolic lattices; see Figure 5.
@@ 5.1. Toroidal compactification for the Coxeter fans
The Coxeter semifan ${\mathfrak{F}}_{\rm cox}$ is the semifan with support $\overline{{\mathcal{C}}}_{\mathbb{Q}}$ whose maximal cones are chambers of $W$, i.e. ${\mathfrak{C}}$ and its $W$-images.
@@
    Type III cones. These are in bijection with elliptic subdiagrams of $G$.
@@
By Lemma 2.8 there is an immersion $j\colon F_{\rm En,2}\to F_{(2,2,0)}$ whose image is a Noether-Lefschetz locus in $F_{(2,2,0)}$. The normalization of the closure of $j(F_{\rm En,2})$ in ${\overline{F}}_{(2,2,0)}^{\rm cox}$ is then a toroidal compactification ${\overline{F}}_{\rm En,2}^{\rm cox}$ for the fans $\{{\mathfrak{F}}^{k}_{\rm cox}\}_{k=1,2,3,4,5}$, one for each of 00-cusp of $F_{\rm En,2}$. The fans ${\mathfrak{F}}^{k}_{\rm cox}$ are the intersections of the above fans ${\mathfrak{F}}_{r}$ in the lattices ${\overline{T}}_{\rm dP}=(18,0,0)_{1}$ and $(18,2,0)_{1}$ with the sublattices ${\overline{T}}_{\rm En}=(10,10,0)_{1}$ and $(10,8,0)_{1}$, as in Section 3. By Lemma 3.8 and Corollary 3.12 these five fans are the Coxeter fans for the folded Coxeter diagrams $G_{r}^{k}$ of Figures 7, 8. By [Ste91] the induced groups acting on $e^{\perp}/e$ are of the form $\Gamma_{k}=\operatorname{Aut}(G_{r}^{k})\ltimes W(G_{r}^{k})$
@@
For $k=1,2,3,4,5$, the numbers of Type II + Type III rays in ${\mathfrak{F}}^{k}_{\rm cox}/\Gamma_{k}$ are $4+4$, $2+8$, $3+15$, $4+12$, $5+17$. The toroidal compactification ${\overline{F}}_{\rm En,2}^{\rm cox}$ has $9+56=65$ Type II + Type III divisors.